solve 7^(x^2)=2^(x+1) for x
Answer by stanbon(57967)

solve 7^(x^2)=2^(x+1) for x
--------
log2(7) = log(7)/log(2) = 2.8074
----
Resrite the problem:
[2^2.8074)^(x^2) = 2^(x+1)
----
2^(2.8074x^2) = 2^(x+1)
----
2.8074x^2 = x+1
-----
2.8074x^2 - x - 1 = 0
----
x = -0.4447... or x = 0.8009...
=================
Cheers,
Stan H.